# Knowledge-guided machine learning model with soil moisture for corn yield prediction under drought conditions

### v1

@@ 2.2.2 Field-level dataset for pretraining
Others Prediction year
Location (Latitude and longitude)
Historical average corn yield (t/ha) (simulated)
@@ 2.3 GEE county-level dataset
In this section, we introduce how to construct a county-level dataset for finetuning from GEE. All the variables in the APSIM field-level dataset were also included. To enhance the ML model with RS data, we included VIs from satellite imagery: Green Chlorophyll Index (GCI), Enhanced Vegetation Index (EVI), Normalized Difference Water Index (NDWI), and Normalized Difference Vegetation Index (NDVI).
@@ 2.3.1 Vegetation indices
The Moderate Resolution Imaging Spectroradiometer (MODIS) dataset (Justice et al., 1998; Schaaf and Wang, 2015) is a vital source of satellite-derived data with 500 m resolution, providing consistent, high-quality observations of the Earth’s surface. MODIS captures data in multiple spectral bands and offers various products like VIs. These indices are essential for monitoring vegetation health, biomass, water content, and chlorophyll levels across different regions and time scales. They help improve yield prediction by indicating vegetation growth and crop health.
The Green Chlorophyll Index (GCI) is used to estimate the chlorophyll content in plants, which directly relates to plant health and productivity. It is calculated using the near-infrared (NIR) and green spectral bands. GCI is valuable for assessing vegetation health, especially for monitoring crop growth and detecting areas with nutrient deficiencies. High chlorophyll content typically indicates a healthy, productive plant, making GCI an essential tool in agricultural management. GCI is defined as:
@@
The Enhanced Vegetation Index (EVI) is designed to optimize the vegetation signal by enhancing sensitivity in high biomass areas and reducing atmospheric and canopy background noise. It uses the red, blue, and NIR spectral bands. EVI is particularly useful in regions with dense vegetation, where NDVI might saturate. It provides a more accurate representation of vegetation conditions, especially in areas with high leaf area index, making it ideal for monitoring forests and agricultural lands. EVI is defined as:
@@
The Normalized Difference Water Index (NDWI) measures the moisture content in vegetation and soil by using the NIR and shortwave infrared (SWIR) bands. NDWI is crucial for monitoring water stress in plants, assessing drought conditions, and detecting changes in water bodies. It aids in managing irrigation practices and understanding the impact of water availability on crop yields. NDWI is defined as:
@@
The Normalized Difference Vegetation Index (NDVI) is one of the most widely used indices for measuring vegetation greenness and biomass. It is calculated using the NIR and red spectral bands. NDVI is a standard tool for monitoring plant growth, assessing vegetation cover, and detecting changes in biomass over time. It is particularly useful for tracking seasonal variations in vegetation and estimating crop yields. NDVI is defined as:
@@ 2.3.2 Weather data
The MCD18A1 Version 6.1 (Wang, 2021) is a MODIS Terra and Aqua combined Downward Shortwave Radiation gridded Level 3 product that provides us with reliable radiation data (Radn). It is produced daily at 500 m resolution, with estimates of Downward Shortwave Radiation provided every 3 hours. Downward Shortwave Radiation is incident solar radiation over land surfaces in the shortwave spectrum (300-4,000 nanometers) and is an important variable in land-surface models that address a variety of scientific and applied issues.
@@ 2.3.3 Soil moisture
The SPL4SMGP.007 SMAP L4 Global 3-hourly 9-km Surface and Root Zone Soil Moisture dataset (Entekhabi et al., 2010) plays a critical role in our research on the relationship between drought and corn yield prediction. By providing detailed measurements of soil moisture at both the surface (0–5 cm) and root zone levels (0–100 cm), SMAP data allows us to assess the availability of water in the soil, a key factor influencing crop growth and resilience during drought conditions.
2.3.4 County-level dataset for finetuning
In addition to the primary weather and vegetation parameters, we also incorporated features such as the prediction year, location (latitude and longitude), and the 5-year historical average yield (USDA, 2020) into our GEE county-level dataset, as in the APSIM field-level dataset. By integrating these features, the model gained a more comprehensive understanding of the environmental and temporal factors affecting corn yield, leading to more reliable predictions. All the variables in GEE county-level dataset are listed in Table 4.
@@ 3 Methodology
The overall pipeline of our KGML-SM is shown in Fig. 2. Our deep learning model consisted of two main components: the W2S encoder, which learned the relationship between weather data and soil moisture, and the attention-based feature-weighting module, which captured how various features influence corn yield. We first pretrained the model using the APSIM field-level dataset, then finetuned it with the GEE county-level dataset.
@@
We begin with a formal Problem formulation of KGML-SM (Sec. 3.1), followed by an introduction to the model components (Sec. 3.2); next, we explain how field-level dataset is generated for pretraining (Sec. 3.3) and county-level dataset for finetuning and testing (Sec. 3.4); finally, we introduce how the KGML model is trained and used for prediction (Sec. 3.5).
@@ 3.1 Problem formulation
We formally define the corn yield prediction problem and present its mathematical formulation within a KGML framework. For a unique combination of county and year sample $i$, where $i\in\mathcal{N}$, $\mathcal{N}$ represents all counties across all years in the experiment, we define the input features as follows. The weather features are denoted as $\mathbf{w_{i}}=[w^{1}_{i},w^{2}_{i},\dots,w^{t}_{i}]$, and the other features (prediction year, location, and historical average yield) as $\mathbf{o_{i}}=[o^{1}_{i},o^{2}_{i},\dots,o^{t}_{i}]$, where $t\in\mathcal{T}$ and $\mathcal{T}$ represents the entire period from corn planting to harvest. We also have the simulated soil moisture as $\mathbf{s_{i}}=[s^{1}_{i},s^{2}_{i},\dots,s^{t}_{i}]$ and $y_{i}$ denotes the corn yield of each county in that year as supervision. The APSIM field-level dataset, $\mathcal{D}_{\text{sim}}=\{(\mathbf{w_{i}},\mathbf{o_{i}}\mid\mathbf{s_{i}},\mathbf{y_{i}})\}$, is used for pretraining. The vertical bar $\mid$ separates the input variables from the target labels.
Additionally, the VIs are represented as $\mathbf{v_{i}}=[v^{1}_{i},v^{2}_{i},\dots,v^{t}_{i}]$. We denote two subsets of the GEE county-level dataset as $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ respectively, where $\mathcal{D}_{\text{train}}$ is used for finetuning and $\mathcal{D}_{\text{test}}$ for testing, $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}=\{(\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}}\mid\mathbf{s_{i}},\mathbf{y_{i}})\}$. Here, we omit the description of the validation set $\mathcal{D}_{\text{val}}$ but introduce its partitioning and usage details in the experimental setup (Sec. 4.1).
The objective is to first learn an encoder $f_{\text{W2S}}$ that maps weather inputs to soil moisture: $\mathbf{\hat{s}_{i}}=f_{\text{W2S}}(\mathbf{w_{i}})$. Then, the predicted soil moisture $\mathbf{\hat{s}_{i}}$ is used along with other input features to predict yield through an attention module $f_{\text{att}}$: $y_{i}=f_{\text{att}}(\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}},\mathbf{\hat{s}_{i}})$. The model’s performance is evaluated by comparing the predicted yields $\hat{y}_{i}$ with the actual yields $y_{i}$.
@@ 3.2.1 Weather-to-Soil encoder
W2S encoder is a module designed to model the influence from weather conditions to soil moisture. By capturing the relationship between weather and soil moisture, the W2S encoder enhances the model’s understanding of soil dynamics and the impact of weather variability on soil conditions and corn yield.
The W2S encoder follows a U-Net-based encoder-decoder structure (Ronneberger et al., 2015), consisting of an encoding function $\mathcal{E}$, a decoding function $\mathcal{D}$, and a final fully connected layer $\mathcal{F}$ for feature transformation. Given an input time-series weather input $\mathbf{w_{i}}$, the encoding process can be formulated as:
$\mathbf{z}_{1}=\mathcal{E}_{1}(\mathbf{w_{i}}),\quad\mathbf{z}_{2}=\mathcal{E}_{2}(\mathcal{P}(\mathbf{z}_{1}))$ (5)
where $\mathcal{E}_{i}$ denotes the $i$-th convolutional block (LeCun et al., 1995), and $\mathcal{P}$ represents the max-pooling operation (Nagi et al., 2011) that reduces the temporal dimension. For simplicity we omit the subscript $i$ for $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$.
The decoding process reconstructs high-level features through upsampling and convolutional layers:
$\mathbf{\tilde{z}}_{2}=\mathcal{U}(\mathbf{z}_{2}),\quad\mathbf{\tilde{z}}_{1}=\mathcal{D}_{1}([\mathbf{\tilde{z}}_{2},\mathbf{z}_{1}])$ (6)
where $\mathcal{U}$ is a transposed convolution (upsampling) operation, $[\mathbf{\hat{z}}_{2},\mathbf{z}_{1}]$ represents skip connections concatenating encoder and decoder feature maps, and $\mathcal{D}_{1}$ is the decoder convolutional block.
The predicted soil moisture $\mathbf{\hat{s}_{i}}$ is obtained via:
$\mathbf{\hat{s}_{i}}=\mathcal{F}(\mathcal{D}_{1}([\mathbf{\tilde{z}}_{2},\mathbf{z}_{1}]))$ (7)
@@ 3.2.2 Attention module
The attention mechanism (Vaswani, 2017) is a powerful tool in ML that enables models to focus on the most relevant parts of the input data when making predictions. By assigning different levels of importance, or weights, to various input elements, the attention mechanism helps the model prioritize the most crucial information for the task. In KGML-SM, we aim to use the attention mechanism to weight different features, helping us understand each feature’s contribution to yield prediction across different dimensions.
Specifically, suppose that we have input $x_{i}=[\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}},\mathbf{\hat{s}_{i}}]$ for predicting corn yield. For each feature embedding $x_{i}$, we use the attention mechanism to learn the corresponding weight $\alpha_{i}$, which is then used to compute the final yield $y_{i}$.
First, we compute the query $Q_{i}$, key $K_{i}$, and value $V_{i}$ vectors from the feature $x_{i}$ using learned linear transformations:
$Q_{i}=W_{Q}x_{i},\quad K_{i}=W_{K}x_{i},\quad V_{i}=W_{V}x_{i}$ (8)
where $W_{Q}$, $W_{K}$, and $W_{V}$ are the learned weight matrices for the query, key, and value, respectively.
@@
$\text{Attention}(Q_{i},K_{i})=\frac{Q_{i}\cdot{K_{i}}^{\top}}{\sqrt{d_{k}}}$ (9)
These attention scores are then passed through a softmax function (Goodfellow et al., 2016) to obtain the weights $\alpha_{i}$:
$\alpha_{i}=\text{softmax}\left(\frac{Q_{i}\cdot{K_{i}}^{\top}}{\sqrt{d_{k}}}\right)$ (10)
And The softmax function (Goodfellow et al., 2016) is defined as:
@@
Where $\mathbf{z}=[z_{1},\dots,z_{K}]$ is the input vector for $i=1,\dots,K$. $e^{z_{i}}$ is the exponential of the $i$-th element of the input vector $\mathbf{z}$.
Finally, we compute the weighted sum of the values $V_{i}$ to obtain the final yield prediction:
$y_{i}=\sum_{t=1}\alpha_{i}V_{i}$ (12)
This process allows us to capture the importance of each feature $x_{i}$ in predicting the yield, dynamically adjusting the weights based on the learned attention scores.
3.3 Generating field-level dataset for pretraining
@@ 3.3.1 APSIM management parameters
In addition to weather input data, the APSIM model requires management parameters for calibration. Simulation methods vary depending on the scale of the study. For field-level crop areas, management information is obtained directly from farmers and used as input, while other unknown parameters are adjusted to optimize the model (Shahhosseini et al., 2021). For county-level crop areas, a common practice is to identify representative fields within those counties, average their management parameters, and apply these averages to represent the entire county (Puntel et al., 2016). When the study area is very large, as in our research on simulating all counties in the U.S. Corn Belt, calibrating each county individually became impractical. In this case, the typical approach is to define a range for each parameter based on empirical and statistical data, and then randomly combine these parameters within the specified ranges to simulate yield across all counties (Lobell et al., 2015). Given the vast area of cornfields in our study, it was not feasible to calibrate these parameters for every individual station. Therefore, we established a general range for these parameters to encompass all possible values across different fields. This approach ensured that the model remained applicable to a wide range of conditions while maintaining a reasonable level of accuracy in its predictions.
We selected an appropriate adjustment range for the most important management parameters (Table 5). Sowing density determines the number of plants per unit area, directly affecting competition for resources such as light, water, and nutrients. Based on statistics from the USDA NASS (USDA, 2020), the range for sowing density was set at 6-9 plants/m2. Sowing dates are crucial as they determine the crop’s growth cycle and its interaction with seasonal weather patterns. Typically, growers maximize corn yield by planting in late April or early May (Licht, 2021; Coulter, 2024). In the simulation, sowing started between April 20–25 and ended between May 15–20. Fertilizer application is essential for providing the necessary nutrients to support plant growth. The most commonly used nitrogen fertilizers for corn production in North America are anhydrous ammonia, urea, and urea-ammonium nitrate solutions (Herzmann et al., 2004). The fertilizer amount was set at 200-300 kg/ha of urea nitrogen (N). Initial soil water content is important for establishing the starting conditions for the model’s simulation of soil moisture dynamics throughout the growing season. The initial soil water content was set between 40% and 60%.
@@
3.3.2 Filtering simulated data based on soil moisture
PB models overlook certain environmental factors and real-world influences, leading to significant discrepancies between some simulated and real-world data. In the simulation process, the simulated intermediate variable is soil moisture, and the final simulation output is corn yield. It is reasonable to assume that with multiple intermediate variables, the final simulation results will have greater errors than the intermediate variables themselves. Additionally, soil moisture is influenced by external factors in a relatively direct manner, unlike yield, which depends on complex plant growth processes. Therefore, soil moisture was used as a benchmark to filter the simulated dataset. We trained a linear regression (LR) model on the GEE county-level dataset to predict soil moisture from weather data and used it to filter the most accurate simulated data in the APSIM field-level dataset. Simulated soil moisture with higher accuracy and its corresponding corn yield were retained to ensure that the model benefited from the most reliable simulations. Notably, we used only simulated data from years before the test year, eliminating the risk of future information leakage.
3.4 Generating county-level dataset for finetuning and testing
3.4.1 Feature Extraction Within Cropland
After downloading various feature data from GEE, we needed to identify the cropland areas. We used the cropland mask from the Cropland Data Layer (CDL) (USDA-NASS, 2017) to extract the feature data corresponding to cropland. The Cropland Data Layer is an annual, raster-based crop-specific land cover dataset produced by the USDA. It has a resolution of 30 meters and offers detailed crop type and location data across the United States. For our corn yield prediction study in twelve U.S. Corn Belt states, we used the CDL to generate a corn field mask in GEE. This mask identifies the specific areas where corn is grown, allowing us to focus our analysis on relevant agricultural fields.

### v2

@@ 2.2.2 Variables summary in field-level dataset for pretraining
Others Prediction year (input)
Location (Latitude and longitude) (input)
Historical average yield (t/ha) (simulated)
@@ 2.3 GEE county-level dataset
In this section, we introduce how to construct a county-level dataset for finetuning. Details of the studied counties are provided in Table 4. This dataset includes all the variable types in the APSIM field-level dataset (Table 3), including weather, soil moisture, and other features, but from different sources. To enhance ML model performance with RS data, we also included four Vegetation Indices (VIs) from satellite imagery: Green Chlorophyll Index (GCI), Enhanced Vegetation Index (EVI), Normalized Difference Water Index (NDWI), and Normalized Difference Vegetation Index (NDVI).
Table 4: The average number of counties selected for each state in GEE county-level dataset.
State IL IN IA KS MI MN MO NE ND OH SD WI
County number 91 74 87 74 57 64 74 69 48 72 39 62
@@ 2.3.1 Vegetation indices
The Terra and Aqua combined Moderate Resolution Imaging Spectroradiometer (MODIS) Land Cover Climate Modeling Grid Version 6 data product (MCD12Q1 v6) (Schaaf and Wang, 2015) is a vital source of satellite-derived data with 500 m resolution, providing consistent, high-quality observations of the Earth’s surface. MODIS MCD12Q1 v6 captures data in multiple spectral bands and offers various products like VIs. These indices are essential for monitoring vegetation health, biomass, water content, and chlorophyll levels across different regions and time scales. They help improve yield prediction by indicating vegetation growth and crop health.
The Green Chlorophyll Index (GCI) (Gitelson et al., 2005) is used to estimate the chlorophyll content in plants, which directly relates to plant health and productivity. It is calculated using the near-infrared (NIR) and green spectral bands. GCI is valuable for assessing vegetation health, especially for monitoring crop growth and detecting areas with nutrient deficiencies. High chlorophyll content typically indicates a healthy, productive plant, making GCI an essential tool in agricultural management. GCI is defined as:
@@
The Enhanced Vegetation Index (EVI) (Huete et al., 2002) is designed to optimize the vegetation signal by enhancing sensitivity in high biomass areas and reducing atmospheric and canopy background noise. It uses the red, blue, and NIR spectral bands. EVI is particularly useful in regions with dense vegetation, where NDVI might saturate. It provides a more accurate representation of vegetation conditions, especially in areas with high leaf area index, making it ideal for monitoring forests and agricultural lands. EVI is defined as:
@@
The Normalized Difference Water Index (NDWI) (Gao, 1996) measures the moisture content in vegetation and soil by using the NIR and shortwave infrared (SWIR) bands. NDWI is crucial for monitoring water stress in plants, assessing drought conditions, and detecting changes in water bodies. It aids in managing irrigation practices and understanding the impact of water availability on crop yields. NDWI is defined as:
@@
The Normalized Difference Vegetation Index (NDVI) (KRIEGLER, 1969) is one of the most widely used indices for measuring vegetation greenness and biomass. It is calculated using the NIR and red spectral bands. NDVI is a standard tool for monitoring plant growth, assessing vegetation cover, and detecting changes in biomass over time. It is particularly useful for tracking seasonal variations in vegetation and estimating crop yields. NDVI is defined as:
@@ 2.3.2 Weather data
The MCD18A1 Version 6.1 (Wang, 2021) is a MODIS Terra and Aqua combined Downward Shortwave Radiation gridded Level 3 product. The reliable radiation data (Radn) is produced daily at 500 m resolution, with estimates of Downward Shortwave Radiation provided every 3 hours. Downward Shortwave Radiation is incident solar radiation over land surfaces in the shortwave spectrum (300-4,000 nanometers) and is an important variable in land-surface models that address a variety of scientific and applied issues.
@@ 2.3.3 Soil moisture
The SPL4SMGP.007 SMAP L4 Global 3-hourly 9-km Surface and Root Zone Soil Moisture dataset (Reichle et al., 2022) plays a critical role in our research on the relationship between drought and corn yield prediction. By providing detailed measurements of soil moisture at both the surface (0–5 cm) and root zone levels (0–100 cm), SMAP level 4 data allows us to assess the availability of water in the soil, a key factor influencing crop growth and resilience during drought conditions.
2.3.4 Variables summary in county-level dataset for finetuning
Additional features included the prediction year, location (latitude and longitude), and the 5-year historical average yield (USDA, 2020). These features were also added to the GEE county-level dataset, consistent with their inclusion in the APSIM field-level dataset. All the variables in GEE county-level dataset are listed in Table 5.
@@ 3 Methodology
The overall pipeline of the KGML-SM is shown in Fig. 2. This architecture comprised two principal components: the W2S encoder, designed to capture the relationship between weather data and soil moisture, and the attention-based feature-weighting module, which learned how various features influence corn yield. Initially, the model was pretrained on the APSIM field-level dataset, followed by finetuning using the GEE county-level dataset.
@@
We begin with a formal problem formulation of KGML-SM (Sec. 3.1), followed by an introduction to the model components (Sec. 3.2); next, we explain how the APSIM field-level dataset is generated for pretraining (Sec. 3.3) and the GEE county-level dataset for finetuning and testing (Sec. 3.4); finally, we introduce how the KGML-SM model is trained and used for prediction (Sec. 3.5).
@@ 3.1 Problem formulation
The corn yield prediction problem is formally defined in this section, along with its mathematical formulation within the KGML framework. Let each unique county-year combination be represented by a sample indexed as $i$ ($i=1,2,\dots,N$), where $N$ is the total number of county-year instances in the experiment. For each sample $i$, the input features are specified as follows: The temporal weather features are denoted by the vector $\mathbf{w_{i}}=[w^{1}_{i},w^{2}_{i},\dots,w^{T}_{i}]$. Other features, including the prediction year, geographical location, and historical average yield, are aggregated into the vector $\mathbf{o_{i}}=[o^{1}_{i},o^{2}_{i},\dots,o^{T}_{i}]$, In these representations, $T$ corresponds to the total number of discrete time steps covering the duration from corn planting to harvest. Additionally, simulated soil moisture data is represented as $\mathbf{s_{i}}=[s^{1}_{i},s^{2}_{i},\dots,s^{T}_{i}]$. The corn yield record for sample $i$ is denoted by $y_{i}$ and serves as the supervision for model training.
The APSIM-generated field-level dataset, specified as $\mathcal{D}_{\text{field/pretrain}}=\{(\mathbf{w_{i}},\mathbf{o_{i}}\mid\mathbf{s_{i}},\mathbf{y_{i}})\}$, is employed for the pretraining phase. In this notation, the vertical bar $\mid$ separates the input variables ($\mathbf{w_{i}},\mathbf{o_{i}}$) from the target labels ($\mathbf{s_{i}},\mathbf{y_{i}}$). Furthermore, VIs represented by the vector $\mathbf{v_{i}}=[v^{1}_{i},v^{2}_{i},\dots,v^{T}_{i}]$ are incorporated. The county-level dataset derived from GEE is denoted as $\mathcal{D}_{\text{county/finetune}}=\{(\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}}\mid\mathbf{s_{i}},\mathbf{y_{i}})\}$ and is utilized for the finetuning.
The objective is to first build the W2S encoder $f_{\text{W2S}}$ to map weather inputs to soil moisture $\mathbf{\hat{s}_{i}}=f_{\text{W2S}}(\mathbf{w_{i}})$. Then, the predicted soil moisture $\mathbf{\hat{s}_{i}}$ is combined with other input features to predict yield via an attention module $f_{\text{att}}$, resulting in the prediction $\hat{y_{i}}=f_{\text{att}}(\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}},\mathbf{\hat{s}_{i}})$. The model’s performance is evaluated by comparing the predicted yields $\hat{y}_{i}$ with the actual yields $y_{i}$.
@@ 3.2.1 Weather-to-Soil encoder
The W2S encoder is a module designed to model the influence of weather conditions on soil moisture. By capturing the relationship between weather and soil moisture, the W2S encoder improves the model’s understanding of soil dynamics and the impact of weather variability on soil conditions and corn yield.
The W2S encoder employs a U-Net-based encoder-decoder architecture, which consists of an encoder, a decoder, and a fully connected layer (Goodfellow et al., 2016) for feature transformation. Given a time-series weather input $\mathbf{w_{i}}$, the encoder extracts hierarchical representations by progressively downsampling the temporal dimension. The decoder then reconstructs high-level features using upsampling and skip connections that integrate information from the encoder. Finally, a fully connected layer transforms the decoded features into the predicted soil moisture output $\mathbf{\hat{s}_{i}}=f_{\text{W2S}}(\mathbf{w_{i}})$.
@@ 3.2.2 Attention module
The attention mechanism (Vaswani, 2017) is a powerful tool in ML that enables models to focus on the most relevant parts of the input data when making predictions. By assigning different levels of importance to various input elements, the attention mechanism helps the model prioritize the most crucial information. In KGML-SM, we aim to use the attention mechanism to weight different features, helping us understand each feature’s contribution to yield prediction across different dimensions.
The input for corn yield prediction is formulated as $\mathbf{X_{i}}=[\mathbf{w_{i}};\mathbf{o_{i}};\mathbf{v_{i}};\mathbf{\hat{s}_{i}}]$ representing the concatenation of feature vectors. For each feature embedding $\mathbf{X_{i}}$, an attention mechanism is employed to learn the corresponding attention weight $\mathbf{\alpha_{i}}$. This weight is subsequently utilized in the computation of the final yield $\hat{y_{i}}$.
First, we compute the query $\mathbf{Q_{i}}$, key $\mathbf{K_{i}}$, and value $\mathbf{V_{i}}$ vectors from the feature $\mathbf{X_{i}}$ using learned linear transformations:
$\mathbf{Q_{i}}=\mathbf{W_{Q}}\mathbf{X_{i}},\quad\mathbf{K_{i}}=\mathbf{W_{K}}\mathbf{X_{i}},\quad\mathbf{V_{i}}=\mathbf{W_{V}}\mathbf{X_{i}}$ (5)
where $\mathbf{W_{Q}}$, $\mathbf{W_{K}}$, and $\mathbf{W_{V}}$ are the learned weight matrices for the query, key, and value, respectively.
@@
$\text{Attention}(\mathbf{Q_{i}},\mathbf{K_{i}})=\frac{\mathbf{Q_{i}}\cdot{\mathbf{K_{i}}}^{\top}}{\sqrt{d_{k}}}$ (6)
These attention scores are then passed through a softmax function (Goodfellow et al., 2016) to obtain the weights $\mathbf{\alpha_{i}}$:
$\mathbf{\alpha_{i}}=\text{softmax}\left(\frac{\mathbf{Q_{i}}\cdot{\mathbf{K_{i}}}^{\top}}{\sqrt{d_{k}}}\right)$ (7)
The softmax function converts a vector of $K$ real numbers into a probability distribution of $K$ possible outcomes. Given a random input vector $\mathbf{z}=[z_{1},\dots,z_{K}]$ for $i=1,\dots,K$, the softmax function (Goodfellow et al., 2016) is defined as:
@@
where $e^{z_{i}}$ is the exponential of the $i$-th element of the input vector $\mathbf{z}$.
Finally, we compute the weighted sum of the values $\mathbf{V_{i}}$ to obtain the final yield prediction:
$\hat{y_{i}}=\sum_{t=1}\mathbf{\alpha_{i}}\mathbf{V_{i}}$ (9)
3.3 Generating APSIM field-level dataset for pretraining
The variables comprising the field-level dataset were previously introduced (Sec. 2.2); this section provides additional processing details. Besides the input weather data obtained from IEM, specific management parameters were also necessary for model calibration (Sec. 3.3.1). Following the simulation, the resulting data required a optimizing process prior to its utilization (Sec. 3.3.2).
@@ 3.3.1 APSIM management parameters
Simulation methodology depends on the study’s scale. For field-level areas, management information is obtained directly from farmers as input. Other unknown parameters are adjusted to optimize the model (Shahhosseini et al., 2021). For county-level areas, identifying representative fields is common practice. Their management parameters are averaged, and then applied to represent the entire county (Puntel et al., 2016). Calibrating each county individually became impractical for very large study areas like our U.S. Corn Belt simulation. In such cases, the typical approach defines a parameter range based on empirical and statistical data. These parameters are then randomly combined within specified ranges, simulating yield across all counties (Lobell et al., 2015). Given the vast area of cornfields in our study, it was not feasible to calibrate these parameters for every individual station. Therefore, we established a general parameter range encompassing all possible values across different fields. This approach ensured model applicability across a wide range of conditions, maintaining reasonable prediction accuracy.
Some reasonable adjustment ranges for key management parameters were collected from prior research and USDA statistics (Table 6). Sowing density determines the number of plants per unit area, directly affecting competition for resources such as light, water, and nutrients. Based on statistics from the USDA NASS (USDA, 2020), the range for sowing density was set at 6-9 plants/m2, values of 6,7,8, and 9 were used in the simulation. Sowing dates are crucial as they determine the crop’s growth cycle and its interaction with seasonal weather patterns. Typically, growers maximize corn yield by planting in late April or early May (Licht, 2021; Coulter, 2024). In the simulation, sowing started between April 20–25 and ended between May 15–20. Fertilizer application is essential for providing the necessary nutrients to support plant growth. The most commonly used nitrogen fertilizers for corn production in North America are anhydrous ammonia, urea, and urea-ammonium nitrate solutions (Herzmann et al., 2004). The fertilizer amount was set at 200-300 kg/ha of urea nitrogen (N), values of 200, 250, and 300 were used in the simulation. Initial soil water content is important for establishing the starting conditions for the model’s simulation of soil moisture dynamics throughout the growing season. The initial soil water content was set between 40% and 60%, values of 40%, 50%, and 60% were used in the simulation.
@@
3.3.2 Optimizing APSIM field-level dataset based on soil moisture
APSIM is not accurate for large-scale corn yield simulation without precise management adjustments, thus requiring optimization of the APSIM field-level pretrain dataset. Soil moisture was proposed as a benchmark for filtering high-quality data. First, soil moisture dynamics are relatively simple, strongly correlated with precipitation. In contrast, corn yield is influenced by multiple factors and complex growth processes, hindering accurate modeling in process-based model. Second, simulation processes inherently exhibit cumulative error. Due to the relative simplicity of soil moisture modeling, this error is comparatively smaller. Therefore, the APSIM field-level dataset was optimized based on the quality of soil moisture data. Specifically, an LR model was trained on the GEE county-level dataset. It utilized four weather inputs (Tmax, Tmin, PPT, Radn) to predict soil moisture output. The trained model subsequently predicted soil moisture for the APSIM field-level dataset using its weather data. This prediction was compared against the simulated soil moisture. Simulated soil moisture with a mean squared error (MSE) exceeding 0.5 were discarded ensuring dataset quality.
3.4 Generating GEE county-level dataset for finetuning and testing
3.4.1 Feature Extraction Within Cornfield
To focus on the corn portion of the acquired feature data from GEE, the cornfield areas had to be identified. The Cropland Data Layer (CDL) cropland mask (USDA-NASS, 2017) was utilized extracting feature data corresponding to cropland. The Cropland Data Layer is an annual raster-based dataset. It provides crop-specific land cover information produced by the USDA. This dataset has a 30-meter resolution. The CDL was employed to generate a cornfield mask in GEE, supporting our corn yield prediction study in twelve U.S. Corn Belt states.